OpenStudy (anonymous):

2x-3y=18 9x+8y=10 FIND X AND Y

OpenStudy (anonymous):

x=174/43 y=-142/43

OpenStudy (anonymous):

from first equation find x in term of y and put in 2nd equation to find y

OpenStudy (isbella):

9x+8y=10 solve for y for the equation 2x-3y=18 y=(-18+2x)/3 plug this in to 9x+8y=10 9x+8((-18+2x)/3)=10 x=173/43 then solve you y value just like felavin answers

OpenStudy (anonymous):

thanks Isbella for this.